# Log-concavity And The Multiplicative Properties of Restricted Partition Functions

### v1

@@ Abstract.
The partition function $p(n)$ and many of its related restricted partition functions have recently been shown independently to satisfy log-concavity: $p(n)^{2}\geq p(n-1)p(n+1)$, and satisfy the inequality: $p(n)p(m)\geq p(n+m)$ with only finitely many instances of equality or failure. This paper proves that this is no coincidence, that any log-concave sequence $\{x_{n}\}$ satisfying a particular initial condition likewise satisfies the inequality $x_{n}x_{m}\geq x_{n+m}$. This paper further determines that these conditions are sufficient but not necessary and considers various examples to illuminate the situation.
@@ 1. Introduction
The partition function $p(n)$ enumerates the number of partitions of a positive integer $n$ where the partitions are positive integer sequences $\lambda=(\lambda_{1},\lambda_{2},...)$ with $0<\lambda_{1}\leq\lambda_{2}\leq~{}\dots$ and $\sum_{j\geq 1}{\lambda_{j}}=n$. For example, $p(4)=5$ since the only ways to partition $4$ are $4$, $3+1$, $2+2$, $2+1+1$, and $1+1+1+1$. Recent interest regarding the various properties of $p(n)$ led to a conjecture by Chen [7], and subsequent proof by DeSalvo & Pak [10], that the sequence $\{p(n)\}_{n\in\mathbb{N}}$ is log-concave for $n>25$. A sequence of integers $\{x_{n}\}_{n\in\mathbb{N}}$ is log-concave if $x_{n}^{2}~{}\geq~{}x_{n-1}x_{n+1}$ and is is log-convex if $x_{n}^{2}\leq x_{n-1}x_{n+1}$ for all $n$.
Log-concavity had been previously established in 1978 by Nicholas [29] for the partition function, but there was a revitalization of interest following the result of DeSalvo & Pak [10]. New results found in the work of Bessenrodt & Ono [30] in 2014 proved that the sequence $\{p(n)\}_{n\in\mathbb{N}}$ satisfies the inequality $p(n)p(m)\geq p(n+m)$ where instances of equality occur at $(2,6),(3,4)$, and instances of failure occur at $(2,2)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,3)$, $(5,5)$ and nowhere else.
For convenience, this paper uses the notation: a sequence $\{x_{n}\}$ is multiplicatively abundant (or simply abundant) if $x_{n}x_{m}\geq x_{n+m}$, and is multiplicatively deficient (or simply deficient) if $x_{n}x_{m}\leq x_{n+m}$ for all but finitely many $n$ and $m$. In these terms, Bessenrodt & Ono [30] proved that $\{p(n)\}$ is abundant. Proofs that the partition function is log-concave or abundant have so far relied on an analytic result of Rademacher type due to Lehmer [26] until 2017 when A. Alanazi, S.M. Gagola III, & A.O. Munagi [2] published a combinatorial proof that $\{p(n)\}$ is abundant.
Many sequences (especially sequences arising from restricted partitions) are sporadic for small indices but become and stay log-concave for sufficiently large indices. The main goal of this paper is to prove that any sequence that is eventually log-concave is also eventually abundant, provided the sequence also satisfies a necessary (but not sufficient) initial condition.
@@
Let $\{x_{n}\}$ be a sequence of positive real numbers and let $N$ be a positive integer such that for all $n>N$, the sequence satisfies $x_{n}^{2}\geq x_{n-1}x_{n+1}$. Then for all $n>N$ and $m\geq 0$, the sequence $\{x_{n}\}$ satisfies:
@@
Moreover, if
$(x_{N+1})^{\frac{1}{N+1}}\geq(x_{N+2})^{\frac{1}{N+2}}$ (2)
then $x_{n+m}\leq x_{n}x_{m}$ for all $n,m>N$.
Note that if $\{x_{n}\}$ is log-concave for all $n\geq 1$, then
$x_{1}^{\frac{m}{n+m-1}}(x_{n+m})^{\frac{n-1}{n+m-1}}\leq x_{n}\leq x_{1}\left(\frac{x_{2}}{x_{1}}\right)^{n-1}$
for all non-negative integer $m$. In light of Theorem 1.1 and because log-concavity has been established for $n>25$ [10], to show the partition function is abundant it suffices to show that Condition (2) holds for $p(26)=2436$ and $p(27)=3010$: it is easy to verify that in fact, $2436^{\frac{1}{26}}>3010^{\frac{1}{27}}$. An example demonstrates that Condition (2) in Theorem 1.1, for a log-concave sequence to be abundant, is required. If a log-concave sequence does not satisfy $(x_{N+1})^{\frac{1}{N+1}}\geq(x_{N+2})^{\frac{1}{N+2}}$, then it is not necessarily true that $x_{n+m}\leq x_{n}x_{m}$ for all $n,m\geq N+1$.
Example 1. Define the sequence $\{a_{n}\}$ by $a_{n}=2^{n+1}-1$ for $n=0,1,2,\ldots$
$a_{n}=1,3,7,15,31,63,127,255,511,1023,2047,4095,8191,16383,32767,65535,\ldots$
The sequence is strictly log-concave:
$12^{2n+2}-2^{n+2}+1=(2^{n+1}-1)^{2}>(2^{n}-1)(2^{n+2}-1)=2^{2n+2}-2^{n}-2^{n+2}+$ 1
However, the sequence $\{(2^{n+1}-1)^{\frac{1}{n+1}}\}$ for $n=0,1,2,\ldots$, is a strictly increasing sequence whose limit equals $2$ as $n$ tends to infinity. Hence, $a(n)$ does not satisfy Condition (2):
$(2^{n+1}-1)^{\frac{1}{n+1}}<(2^{n+2}-1)^{\frac{1}{n+2}}.$
In this particular case, $\{a_{n}\}$ is also not abundant (in fact, $\{a_{n}\}$ is deficient):
$(2^{n+1}-1)(2^{m+1}-1)=2^{n+m+2}-2^{n}-2^{m}+1<2^{n+m}-1$
for $n,m\geq 1$.
The interaction between log-behavior and multiplicative behavior in the setting of general sequences was been studied in the early 2000’s by Asai, Kubo, & Kuo [3] and by Došlić [12] who collectively established inequalities for Bell numbers, Motzkin numbers, Fine numbers, Fresnel numbers, Apèry numbers, large Schröder numbers, central Delannoy numbers, and Baxter permutations. Both Asai, Kubo, & Kuo [3] and Došlić [12] give a proof that any log-convex sequence is deficient, provided $x_{0}=1$. More recently, a similar version of this proof emerged in 2023 by Gajdzica, Miska, & Ulas [17] with the slight modification that $x_{0}\geq 1$. Unfortunately, there is some misunderstanding that every log-convex sequence is deficient. Note that the reciprocal of a log-concave sequence of positive numbers is a log-convex sequence, and vice versa.
$x_{n}^{2}\geq x_{n-1}x_{n+1}\qquad\iff\qquad\left(\frac{1}{x_{n}}\right)^{2}\leq\left(\frac{1}{x_{n-1}}\right)\left(\frac{1}{x_{n+1}}\right)$
Then the reciprocal sequence in Example 2 is a counterexample to Theorem 2(a) of Asai, Kubo, & Kuo [3] and to Proposition 2.1(b) of Došlić [12] and to Theorem 4.4 of Gajdzica, Miska, & Ulas [17].
@@
Example 2. Define the sequence $\{b_{n}\}$ as the reciprocal of $a_{n}=2^{n+1}-1$ (as in Example 1) for $n=0,1,2,\ldots$ Then, because $a_{n}$ is strictly log-concave, the reciprocal sequence $b_{n}$ is strictly log-convex: $b_{n}^{2}>b_{n-1}b_{n+1}$. However, the sequence does not satisfy Condition (3):
$(2^{n+1}-1)^{\frac{1}{n+1}}<(2^{n+2}-1)^{\frac{1}{n+2}}\qquad\iff\qquad(2^{n+1}-1)^{\frac{-1}{n+1}}>(2^{n+2}-1)^{\frac{-1}{n+2}}.$
Further, $b_{n}$ is not deficient (in fact, $b_{n}$ is abundant):
$(2^{n+1}-1)^{-1}(2^{m+1}-1)^{-1}>(2^{n+m}-1)^{-1}$
for $n,m\geq 1$.
@@
Moreover, if
$(y_{N+1})^{\frac{1}{N+1}}\leq(y_{N+2})^{\frac{1}{N+2}},$ (3)
then $y_{n+m}\geq y_{n}y_{m}$ for all $n,m>N$.
Again, note that if $\{y_{n}\}$ is log-convex for all $n\geq 1$, then
$y_{1}^{\frac{m}{n+m-1}}(y_{n+m})^{\frac{n-1}{n+m-1}}\geq y_{n}\geq y_{1}\left(\frac{y_{2}}{y_{1}}\right)^{n-1}$
for all $n$ and $m$. Example 2 demonstrates that Condition (3) in Theorem 1.2, for a log-convex sequence to be deficient, is required. If a log-convex sequence does not satisfy $(y_{N+1})^{\frac{1}{N+1}}\leq(y_{N+2})^{\frac{1}{N+2}}$, then it is not necessarily true that $x_{n+m}\geq x_{n}x_{m}$ for all $n,m\geq N+1$.
@@ 1.1. Sufficient but not Necessary
$a_{n}=\begin{cases}2,&\text{if}\ n\equiv 0\mod{4}\\ 3,&\text{if}\ n\equiv 1,3\mod{4}\\ 4,&\text{if}\ n\equiv 2\mod{4}\end{cases}$
$a_{n}=2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,\ldots\qquad\text{for}\ n=0,1,2,\ldots$
The inequality $a_{n}^{2}\geq a_{n-1}a_{n+1}$ fails whenever $n\equiv 0\mod{4}$, and so $\{a_{n}\}$ is never log-concave. Similarly, the inequality $a_{n}^{2}\leq a_{n-1}a_{n+1}$ fails whenever $n\equiv 2\mod{4}$, and so $\{a_{n}\}$ is never log-convex. However, the sequence $\{a_{n}\}$ satisfies the inequality $a_{n}a_{m}\geq a_{n+m}$ for all $n$ and $m$. Note that $(a_{n+1})^{\frac{1}{n+1}}\leq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 0,3\mod{4}$ and $(a_{n+1})^{\frac{1}{n+1}}\geq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 1,2\mod{4}$.

### v4

@@ Abstract.
The partition function $p(n)$ and many of its related restricted partition functions have recently been shown independently to satisfy log-concavity: $p(n)^{2}\geq p(n-1)p(n+1)$ for $n\geq 26$, and satisfy the inequality: $p(n)p(m)\geq p(n+m)$ for $n\geq m\geq 2$ with only finitely many instances of equality or failure. This paper proves that this is no coincidence, that any log-concave sequence $\{x_{n}\}$ satisfying a particular initial condition likewise satisfies the inequality $x_{n}x_{m}\geq x_{n+m}$. This paper further determines that these conditions are sufficient but not necessary and considers various examples to illuminate the situation.
@@ 1. Introduction
The partition function $p(n)$ enumerates the number of partitions of a positive integer $n$ where the partitions are positive integer sequences $\lambda=(\lambda_{1},\lambda_{2},...)$ with $0<\lambda_{1}\leq\lambda_{2}\leq~{}\dots$ and $\sum_{j\geq 1}{\lambda_{j}}=n$. For example, $p(4)=5$ since the only ways to partition $4$ are $4$, $3+1$, $2+2$, $2+1+1$, and $1+1+1+1$. Recent interest regarding the various properties of $p(n)$ led to a conjecture by Chen [9], and subsequent proof by DeSalvo & Pak [12], that the sequence $\{p(n)\}_{n\in\mathbb{N}}$ is log-concave for $n>25$. A sequence of integers $\{x_{n}\}_{n\in\mathbb{N}}$ is log-concave if $x_{n}^{2}~{}\geq~{}x_{n-1}x_{n+1}$ and is log-convex if $x_{n}^{2}\leq x_{n-1}x_{n+1}$ for all $n\geq n_{0}$ for some $n_{0}\in\mathbb{N}$.
Log-concavity had been previously established in 1978 by Nicholas [32] for the partition function, but there was a revitalization of interest following the result of DeSalvo & Pak [12]. New results found in the work of Bessenrodt & Ono [6] in 2014 proved that the sequence$\{p(n)\}_{n\in\mathbb{N}}$ satisfies the inequality
$p(n)p(m)\geq p(n+m)\quad\text{for all}\quad(m,n)\quad\text{with}\quad n\geq m\geq 2$ (1)
where instances of equality occur at $(2,6),(2,7),(3,4)$, and instances of failure occur at $(2,2)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,3)$, $(3,5)$ and nowhere else.
For convenience, this paper uses the notation: a sequence $\{x_{n}\}$ is multiplicatively abundant (or simply abundant) if $x_{n}x_{m}\geq x_{n+m}$, and is multiplicatively deficient (or simply deficient) if $x_{n}x_{m}\leq x_{n+m}$ for all $n\geq m\geq n_{0}$ but finitely many $n\geq m\geq n_{0}$. In these terms, Bessenrodt & Ono [6] proved that $\{p(n)\}$ is abundant for all $n\geq m\geq 2$. Proofs that the partition function is log-concave or abundant initially relied on an analytic result of Rademacher type due to Lehmer [28] until 2017 when A. Alanazi, S.M. Gagola III, & A.O. Munagi [2] published a combinatorial proof that $\{p(n)\}$ is abundant.
Many sequences (especially sequences arising from restricted partitions) are sporadic for small indices but become and stay log-concave for sufficiently large indices. The interaction between log-behavior and multiplicative behavior in the setting of general sequences was studied in the early 2000’s by Asai, Kubo, & Kuo [3] who collectively established inequalities for Bell numbers, Motzkin numbers, Fine numbers, Fresnel numbers, Apèry numbers, large Schröder numbers, central Delannoy numbers, and Baxter permutations. Recently, Gajdzica, Miska, & Ulas [18] studied similar interactions in a more general setting.
The main goal of this paper is to prove that any sequence that is eventually log-concave is also eventually abundant, provided the sequence also satisfies a sufficient initial conditions.
@@
Let $\{x_{n}\}$ be a sequence of positive real numbers and let $N$ be a non-negative integer such that for all $n>N$, the sequence satisfies $x_{n}^{2}\geq x_{n-1}x_{n+1}$. Then for all $n>N$ and $m\geq 0$, the sequence $\{x_{n}\}$ satisfies:
@@
Moreover, if there exists an integer $k\geq 0$ such that
$(x_{N+k})^{N+k+1}\geq(x_{N+k+1})^{N+k}$ (3)
then $x_{n+m}\leq x_{n}x_{m}$ for all $n,m\geq N+k$. In addition, $\{x_{n}\}$ is abundant for all $m,n\geq d$ with at most finite number of failures $(m,n)\in[d,M+1]\times[d,N+k]\bigcup[d,N+k]\times[d,M+1]$, if $d\leq N+k$ is the smallest positive integer such that
$\left(\frac{x_{M+1}}{x_{M}}\right)^{m}\leq x_{m}$ (4)
for all $d\leq m\leq N+k$ and for some $M\geq N+k$.
@@
Remarks: 1) If a log-concave sequence does not satisfy condition (3), i.e., $(x_{N+k})^{N+k+1}\geq(x_{N+k+1})^{N+k}$ for all $k\geq 0$, then it is not true that $x_{n+m}\leq x_{n}x_{m}$. For example, define the sequence $\{a_{n}\}$ by $a_{n}=2^{n}-1/2$ for $n=0,1,2,\ldots$ The sequence is strictly log-concave:
$\left(2^{n}-\frac{1}{2}\right)^{2}=2^{2n}-2^{n}+\frac{1}{4}>2^{2n}-2^{n}-2^{n-2}+\frac{1}{4}=\left(2^{n+1}-\frac{1}{2}\right)\left(2^{n-1}-\frac{1}{2}\right).$
However, the sequence $\{(2^{n}-1/2)^{\frac{1}{n}}\}$ for $n=1,2,3,\ldots$, is a strictly increasing sequence. Hence, $a(n)$ does not satisfy Condition (3). Indeed, $(2^{n+k}-1/2)^{n+k+1}<(2^{n+k+1}-1/2)^{n+k}$ for all $n,k\geq 0$. In this particular case, $\{a_{n}\}$ is also not abundant (in fact, $\{a_{n}\}$ is deficient):
$(2^{n}-1/2)(2^{m}-1/2)=2^{n+m}-2^{n-1}-2^{m-1}+\frac{1}{4}<2^{n+m}-1/2$
for $n\geq m\geq 0$.
2) Suppose $N=0$, i.e., $\{x_{n}\}$ is log-concave for all $n\geq 1$, then
$x_{0}^{\frac{m}{n+m}}(x_{n+m})^{\frac{n}{n+m}}\leq x_{n}\leq x_{0}\left(\frac{x_{1}}{x_{0}}\right)^{n}$
for all non-negative integers $m$. This give an upper bound for a log-concave sequence $\{x_{n}\}$. For example, a sequence with super exponential growth can not be log-concave. Moreover, $x_{n+m}\leq x_{n}x_{m}$ for all $n\geq m\geq k$ if $(x_{k})^{k+1}\geq(x_{k+1})^{k}$ for any fixed non-negative integer $k$. Specifically, if $k=0$, meaning that $x_{0}\geq 1$, then for all $n\geq m\geq 0$, $x_{n+m}\leq x_{n}x_{m}$. In 2001, Asai, Kubo, & Kuo [3, Theorem 2] [^1] While $x_{0}=1$ is mentioned in the theorem’s statement, the proof works for $x_{0}\leq 1$. To obtain the proof for the log-concave sequence, one must take the reciprocal of the log-convex sequence in [3]. proved the case $N=0$ and $k=0$ of Theorem 1.1. Gajdzica, Miska, & Ulas [18, Theorem 4.4] proved this case again in $2023$ using a similar technique used in [3]. It can be seen that Theorem 1.1 not only generalized the results of [3, Theorem 2] or [18, Theorem 4.4], it is also more effective. For instance, Bessenrodt-Ono result (1) can be obtained from Theorem 1.1 but it can not be derived from the $N=k=0$ case of Theorem 1.1. Instead one may obtain a weaker result such as $\{p(n+26)\}_{n\geq 0}$ is abundant (see [18, Example 4.5]). A weaker version of Theorem 1.1 for $k=0$ and $M=N$ was recently derived by Heim and Neuhauser [22]. As we shall demonstrate in section 1.1, the result given in [22] does not imply that the log-concave sequences in example 5 are abundant for smaller indices.
1.1. Some Examples of Theorem 1.1
1) In light of Theorem 1.1 and because log-concavity has been established for $n>25$ in [12, 32], to show the partition function is abundant it suffices to show that Condition (3) holds for $p(25)=1958$ and $p(26)=2436$ with $N=25$ and $k=0$: it is easy to verify that in fact, $1958^{\frac{1}{25}}>2436^{\frac{1}{26}}$. It can be noted that this result only proves the partition function is abundant for $n\geq m\geq 25$ but not the result of Bessenrodt-Ono given in (1). To show (1) holds one needs to check condition (4), in particular, for $M=25$: $p(26)/p(25)\leq(p(m))^{1/m}$ for all $2\leq m\leq 25$. This can be easily verified. Next, by checking in the region $[2,26]\times[2,26]$, we see failure occur at $(2,2)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,3)$, $(3,5)$, $(3,2)$, $(4,2)$, $(5,2)$, $(5,3)$ and equality occur at $(2,6),(2,7),(3,4),(6,2),(7,2),(4,3)$. This gives the result of Bessenrodt-Ono (1).
2) The overpartition function, $\overline{p}(n)$, counts the number of partitions of $n$ in which the last occurrence of a number can be distinguished. For example, $\overline{p}(4)=14$:
[…]
Combining the result of Engel [14] and the fact $\overline{p}^{2}(1)\geq\overline{p}(0)\overline{p}(2)$ we see that $\{\overline{p}(n)\}$ is log-concave for $n>0$. Note that $\overline{p}$ satisfies both Conditions (3) and (4) for $N=k=0$. Then Theorem 1.1 give us that the overpartition function satisfies the inequality
$\overline{p}(m)\overline{p}(n)\geq\overline{p}(m+n)\quad\text{forall}\quad m,n\geq 0$ (5)
where instances of equality occur at $(1,1)$, $(1,2)$, $(2,1)$, $(m,0)$ and $(0,n)$. Note that (5) was proven using the asymptotic of the overpartition function by Liu & Zhang [29] in 2021.
3) The distinct Partition Function $P_{d}(n)$ counts the number of ways to partition $n$ into distinct parts, where for all $i$ and $j$, $\lambda_{i}\neq\lambda_{j}$.
[…]
Curiously, Euler showed that $P_{d}(n)$ is equivalent to the number of partitions of $n$ using only odd integers [15, 27, 31]. For example, $P_{d}(4)=2$ since there are only two ways to partition $4$ into distinct parts: $4=4=3+1$ (and only two ways to partition $4$ into odd parts: $4=3+1=1+1+1+1$). In 2023, Dong & Ji [13] proved that the sequence $\{P_{d}(n)\}$ is log-concave for all $n>32$. Note that $P_{d}(n)$ satisfies Condition (3) for $N=32$ and $k=0$: $390^{\frac{1}{32}}>448^{\frac{1}{33}}$ and condition (4) for $M=N=32$: $448/390\leq p_{d}(m)^{1/m}$ for all $3\leq m\leq 32$. Then from Theorem 1.1 and numerical computation gives the distinct partition function satisfies the inequality
$P_{d}(m)P_{d}(n)\geq P_{d}(m+n)\quad\text{for all}\quad n\geq n>2$ (6)
where instances of equality and failure occur as given in Table 1.
Table 1. Equality and Failure for $P_{d}(m)P_{d}(n)\geq P_{d}(m+n)$
Equality at $(m,n)$ Failure at $(m,n)$
$(1,n)$; $n=1,3$ $(1,n),(2,n)$ ; $n\geq 2$
$(3,n)$; $n=3,5,6,7,8$ $(1,2)$, $(3,4)$, $(5,5)$
$(4,n)$; $n=15,16,17$ $(4,n)$; $4\leq n\leq 14,$
$(5,n)$; $n=6,7,8$
The inequality (6) together with the table 1 was proven by Beckwith & Bessenrodt in [4] using a Rademacher type asymptotic formula of $k$-regular partition.
4) This is an example of $k>0$ case. Define the sequences $a_{0}=\frac{1}{2}$ and $a_{n}=2^{\sqrt{n}}$ for $n\geq 1$. Keep in mind that $a_{1}^{2}=2^{2}>\frac{1}{2}\times 2^{\sqrt{2}}=a_{0}a_{2}$, and that $a_{n+1}^{2}=2^{2\sqrt{n+1}}>2^{\sqrt{n}}\times 2^{\sqrt{n+2}}=a_{n}a_{n+2}$ for all $n\geq 1$ and that $a_{n+m}=2^{\sqrt{n+m}}<2^{\sqrt{n}}\times 2^{\sqrt{m}}=a_{n}a_{m}$ for all $n,m\geq 1$. In other words, $\{a_{n}\}$ fulfills the inequality $a_{n+m}<a_{n}a_{m}$ for all $n,m\geq 1$, but is a log-concave sequence for any $n$ with $a_{0}<1$. Since $a_{0}<1$, the result in [18, Theorem 4.4] is not applicable here. It is hopefully clear that $(a_{1})^{2}=2^{2}>2^{\sqrt{2}}=(a_{2})^{1}$ which satisfy Condition (3) in Theorem 1.1.
5) Lastly, we give an example when $M>N+k$. Consider the sequence $a_{n}=\left(\frac{31}{25}\right)^{n}+\frac{1}{10^{6}}$ when $0\leq n\leq 24$ and $a_{n}=p(n)$ when $n\geq 25$. Here $p(n)$ is the partition function. It can be checked that $a_{n}^{2}\geq a_{n-1}a_{n-2}$ for $n\geq 25$. Also, it can be verified that $(a_{n})^{1/n}<\frac{a_{26}}{a_{25}}<\frac{a_{25}}{a_{24}}$ for all $1\leq n\leq 24$ but $(a_{n})^{1/n}>\frac{a_{27}}{a_{26}}$ for all $1\leq n\leq 25$. Thus, using Theorem 1.1, we can deduce that, for every $m,n\geq 1$, $a_{n+m}\leq a_{n}a_{m}$, with the exception of finitely many failures in $[1,25]\times[1,25]$. Note that, the result in [22] yields a weaker result, which only gives $a_{n+m}\leq a_{n}a_{m}$ for all $m,n\geq 24$.
@@
Moreover, if there exists an integer $k\geq 0$ such that
$(y_{N+k})^{N+k+1}\leq(y_{N+k+1})^{N+k},$ (7)
then $y_{n+m}\geq y_{n}y_{m}$ for all $n,m\geq N+k$. In addition, $\{y_{n}\}$ is deficient for all $m,n\geq d$ with at most finite number of failures $(m,n)\in[d,M+1]\times[d,N+k]\bigcup[d,N+k]\times[d,M+1]$, if $d\leq N+k$ is the smallest positive integer such that
$\left(\frac{y_{M+1}}{y_{M}}\right)^{m}\geq y_{m}$ (8)
for all $d\leq m\leq N+k$ and for some $M\geq N+k$.
@@ 1.2. Sufficient but not Necessary
$a_{n}=\begin{cases}2,&\text{if}\ n\equiv 0\mod{4}\\ 3,&\text{if}\ n\equiv 1,3\mod{4}\\ 4,&\text{if}\ n\equiv 2\mod{4}.\end{cases}$
The inequality $a_{n}^{2}\geq a_{n-1}a_{n+1}$ fails whenever $n\equiv 0\mod{4}$, and so $\{a_{n}\}$ is never log-concave. Similarly, the inequality $a_{n}^{2}\leq a_{n-1}a_{n+1}$ fails whenever $n\equiv 2\mod{4}$, and so $\{a_{n}\}$ is never log-convex. However, the sequence $\{a_{n}\}$ satisfies the inequality $a_{n}a_{m}\geq a_{n+m}$ for all $n$ and $m$. For $n>0$, note that $(a_{n+1})^{\frac{1}{n+1}}\leq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 0,3\mod{4}$ and $(a_{n+1})^{\frac{1}{n+1}}\geq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 1,2\mod{4}$.